Please help me free tutor《1,8,27,64》what is the next term
Answer by Theo(13342)

1^3 = 1
2^3 = 8
3^3 = 27
4^3 = 64

the formula is taking the term number and then cubing it.

the first term is term 1.
the second term is term 2.
the third term is term 3.
the fourth term is term 4.

the next term would be term 5, and the value would be 5^3 = 125.